OpenStudy (anonymous):

integral of 3/x from 2 to 4 by hand

OpenStudy (unklerhaukus):

\[\int\limits_2^4\frac 3x\text dx\]

OpenStudy (anonymous):

\[\int\limits_{2}^{4}\frac{ 3 }{ x }dx\]

OpenStudy (unklerhaukus):

the three can be factored out of the integral

OpenStudy (anonymous):

okay, then what? its 3 * (1/x)

OpenStudy (unklerhaukus):

\[\int\limits_a^b\frac 1x\text dx=\left.\ln(x)\right|_a^b=\ln(b)-\ln(a)\]

OpenStudy (unklerhaukus):

can you see what to do now @liodeisel007 ?

OpenStudy (anonymous):

I was working it out and i got to ln(4)-ln(2)

OpenStudy (unklerhaukus):

what happened to the three/?

OpenStudy (anonymous):

its in the front?

OpenStudy (anonymous):

so \[3(2\ln(2)- \ln(2))\] ?

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

thats is right but you can simplify further

OpenStudy (anonymous):

\[3\ln(2)\] ?

OpenStudy (unklerhaukus):

you got it!!!

OpenStudy (anonymous):

thanks!!!